OpenStudy (anonymous):

Integrate by substitution e^(y^(1/2))/y^(1/2) Thanks! Answer is 2e^(y^(1/2)) +C I just can't get there

jimthompson5910 (jim_thompson5910):

Hint: Let u = y^(1/2) so du/dx = (1/2)y^(-1/2) ---> du = 1/(2y^(1/2)) ---> 2*du = 1/(y^(1/2)) dx

jimthompson5910 (jim_thompson5910):

so int( e^(y^(1/2))/y^(1/2) dy ) turns into int( 2*e^u*du ) oh btw, i meant to say dy not dx

OpenStudy (anonymous):

which y^(1/2) is u representing the one in the denominator?